### v1

# NeurDB: On the Design and Implementation of an AI-powered Autonomous Database

Abstract.
Databases are increasingly embracing AI to provide autonomous system optimization and intelligent in-database analytics, aiming to relieve end-user burdens across various industry sectors. Nonetheless, most existing approaches fail to account for the dynamic nature of databases, which renders them ineffective for real-world applications characterized by evolving data and workloads. This paper introduces NeurDB, an AI-powered autonomous database that deepens the fusion of AI and databases with adaptability to data and workload drift. NeurDB establishes a new in-database AI ecosystem that seamlessly integrates AI workflows within the database. This integration enables efficient and effective in-database AI analytics and fast-adaptive learned system components. Empirical evaluations demonstrate that NeurDB substantially outperforms existing solutions in managing AI analytics tasks, with the proposed learned components more effectively handling environmental dynamism than state-of-the-art approaches.
1. Introduction
Database management systems (DBMSs) are becoming more automated and intelligent by embracing artificial intelligence (AI). AI empowers these systems to achieve autonomous optimization Kraska et al. (2019); Pavlo et al. (2017); Zhu et al. (2024) and support complex analytics Karanasos et al. (2020); Xing et al. (2024); Ooi et al. (2015) with minimal human intervention, thereby catering to advanced data- and AI-centric applications.
The aspiration of integrating DBMSs with AI was first expressed forty years ago Brodie (1988), which has been periodically revisited with evolving technology Wang et al. (2016); Cong et al. (2024); Ré et al. (2015). With the advancements in both AI and DBMSs, considerable progress has been made in deepening their fusion. Unfortunately, a huge gap remains between the potential of this integration and its current state of usability, largely due to the inherent differences in their paradigms. The dynamic nature of databases, characterized by constantly changing data and workloads due to transactional updates, and user interactions, poses a fundamental challenge. For example, an e-commerce database might experience a sharp increase in workloads during flash sales compared to normal periods, caused by huge volumes of updates from sales transactions and processing. In contrast, AI models typically derive intelligence from static datasets and thus can become outdated quickly in the face of database dynamism. For example, models for learned query optimization often acquire knowledge by training on specific system environments including workload patterns and data distributions. However, when these environments drift, e.g., due to data updates from user interactions, models relying on past knowledge may lead to ineffective query optimizations since they are not updated to reflect new conditions.
The discrepancy in dynamicity between AI and DBMSs poses a great challenge to the effective adoption of AI-enhanced DBMSs. First, traditional AI-driven system optimization, including knob tuning Pavlo et al. (2017), intelligent advisors Zhu et al. (2024); Li et al. (2021), and learned system components Kraska et al. (2019) such as query optimizers Marcus et al. (2021); Zhu et al. (2023), indexes Kraska et al. (2018), and concurrency control Wang et al. (2021), often overlook the dynamicity of data distributions and workloads. Second, another important facet of AI and DBMS integration, namely in-database AI analytics for complex tasks such as disease progression predictions and user purchase recommendations, has gained traction but faces similar adaptability issues. Existing systems generally focus on static model inference Salazar-Díaz et al. (2024); Park et al. (2022), model serving Vartak (2017), and others Xu et al. (2022), without considering necessary model adaptations to evolving data and workload patterns. Consequently, they require frequent manual updates and retraining to remain effective, which complicates the analytics workflow. Recent works Wu and Ives (2024); Li et al. (2022, 2023a); Kurmanji and Triantafillou (2023) aim to address these challenges but are typically more effective in controlled scenarios, such as data distribution changes following certain rules or query optimization under read-only workloads. These methods have not been designed to generalize across more dynamic and practical workloads involving continuous read and write operations.
We envision a deep integration of AI and DBMSs for adaptability to both data and workload drift. Unlike existing approaches that merely overlay AI onto DBMSs or selectively enhance certain system components, we aim for a comprehensive fusion of AI with DBMSs that enables continuous adaptation to drift. However, achieving this level of integration presents significant challenges. First, the integration necessitates a fundamental redesign of the entire AI workflow, including training, inference, and fine-tuning, within the database architecture. Second, models for learned database components and AI analytics must swiftly adapt to drifting data workloads without losing effectiveness and efficiency. Therefore, developing such a system requires both a system foundation that seamlessly supports model adaptation and AI models that are inherently adaptive in structure.
In this paper, we present NeurDB [^2] https://github.com/neurdb/neurdb, an AI-powered autonomous DBMS that provides efficient and effective in-database AI analytics to seamlessly support modern AI-powered applications, and fully embraces AI techniques in each major system component to offer self-driving capabilities. At the core of NeurDB is an in-database AI ecosystem that achieves deep integration of the AI workflow into the database. In this ecosystem, we develop multiple in-database AI operations, such as model training, inference, and fine-tuning, along with an in-database AI engine to handle the execution of these operators. NeurDB can then directly support in-database AI analytics by calling these AI operations. To further simplify end-users in submitting their complex AI analytics tasks to NeurDB, we provide a user-friendly interface by extending SQL syntax with PREDICT syntax. Moreover, we propose two techniques to optimize the AI ecosystem performance. First, we devise a data streaming protocol that significantly reduces data transfer overhead, yielding better performance for AI operations. Second, we develop an incremental update technique to minimize the fine-tuning cost, facilitating the fast adaptation of AI models. Based on this foundation, we propose two efficient learned system components, namely a learned concurrency control algorithm and a learned query optimizer, that can adapt quickly to data and workload drift.
2. Preliminaries
In this section, we outline the design goals of NeurDB and present the SQL syntax enhanced for AI-powered Analytics.
2.1. Design Goals
DBMSs are dynamic as data distributions and workloads evolve over time, and therefore, the system must be designed for adaptability while also guaranteeing reliability and scalability. Building upon this understanding, we define three essential properties for NeurDB.
Adaptability is the capability of a DBMS to evolve autonomously in response to drifting data and workloads. With optimal adaptability, DBMS can respond to drift in real time.
Reliability depicts the ability of a DBMS to consistently meet performance and accuracy standards, even during phases of adaptation and evolution. With optimal reliability, the system can operate at peak performance and maintain high accuracy consistently.
Scalability refers to the system’s capacity to maintain or enhance performance as the workload increases by introducing more resources, such as threads or nodes.
Informally, the design goal of NeurDB is to ensure the system can uphold reliability as quickly as possible when the system or its components adapt or evolve due to data and workload drift, while ensuring good scalability.
2.2. SQL Syntax for AI Analytics
NeurDB incorporates enhanced SQL, which extends from the standard SQL to support AI analytics. As illustrated in Listing 1 and Listing 2, NeurDB introduces a new PREDICT keyword to handle two typical ML tasks: classification with the CLASS OF clause and regression with the VALUE OF clause. Inspired by the original principle of introducing SQL that allows an application developer to write a query with SELECT, and then let the DBMS find the most efficient way to store and retrieve data, we ensure that a developer can submit an AI analytic task simply with PREDICT. All the following operations, such as retrieving training data and invoking AI models, are handled automatically by the database. We also plan to expose more enhanced SQL for AI model management and other AI analytics functionalities. Due to space constraints, we mainly focus on the PREDICT query in this paper.
We now present two real-world analytic scenarios that can be directly supported by NeurDB using PREDICT.
Classification
The disease progression prediction in healthcare can be handed by the SQL shown in Listing 1. By specifying the features with TRAIN ON and using data stored in the ‘diabetes’ table and directly input with VALUES, we predict the class of the ‘outcome’ variable, which indicates whether a patient has diabetes or not.
Regression
Listing 2 shows the SQL to predict and fill in the missing scores of products based on user reviews. Similarly, it specifies the features for training and the data table to predict the value of the ‘score’ variable. Notably, it automatically excludes features with unique constraints to avoid including meaningless data.
Listing 1: Syntax for a classification task
1PREDICT CLASS OF outcome
2FROM diabetes
3TRAIN ON pregnancies, glucose, blood_pressure, ...
4VALUES (6, 148, 72, ...), (1, 85, 66, ...), ...
Listing 2: Syntax for a regression task
1PREDICT VALUE OF score
2FROM review
3WHERE brand_name = ’SpecialGoods’
4TRAIN ON * WITH brand_name <> ’SpecialGoods’
3. System Overview
Figure 1. The System Architecture of NeurDB
We now describe the system architecture of NeurDB as shown in Figure 1, where we achieve the deep fusion of AI and databases by establishing an in-database AI ecosystem. To achieve this, we first holistically redesign each existing database component, such as the query optimizer, transaction manager, and index manager, to support in-database AI analytics and function as learned components for improved performance. In particular, we enhance the SQL parser and the query optimizer to produce a customized query plan for PREDICT queries. Beyond traditional operators for fetching and processing data, e.g., scan and join, the query executor in NeurDB includes in-database AI operators for model training, inference, and fine-tuning. With these AI operators, we effectively align the AI workflow with the database query processing. We are developing additional AI operators, such as model selection, to provide more comprehensive AI services. For example, a query may call the model selection operator (denoted as MSelection) to automatically select the best-suited model for a given prediction task Xing et al. (2024), thereby enhancing accuracy and efficiency. We then introduce a new component, called the AI engine, into NeurDB. It handles the AI-related processing requests from AI operators and learned components and creates AI tasks on their behalf, where the task manager dispatches them to the CPU/GPU runtime for execution. Specifically, we propose a data streaming protocol to reduce data transfer overhead, and an incremental model update technique to facilitate fast model adaptation, thus enhancing the performance of the AI engine. We shall elaborate on these optimizations in Section 4.1. Also, we design dedicated AI model storage to store AI models and serve them based on requests from the AI engine. We further implement a monitor to detect unexpected performance or accuracy issues, based on which we trigger automatic and appropriate model adaptation.
In Figure 1, we illustrate a running example of a PREDICT query. After parsing and optimizing the query, the executor performs the scan operation to retrieve data and then invokes the inference operator to deliver the inference task to the AI engine. Next, the AI engine uploads the model from the model storage if it is not in the model buffer, and conducts model inference to produce the results.
4. The Design of NeurDB
In this section, we present two key subsystems of NeurDB, namely an in-database AI ecosystem and learned system components.
4.1. In-database AI Ecosystem
Figure 2. AI Engine of NeurDB
AI Engine
The AI engine of NeurDB, pivotal to all AI-related activities for tasks from both user and internal learned components, operates on a distributed and event-driven architecture to optimize efficiency and throughput. Figure 2 shows the main components and the communicative flow with connected external nodes serving distributed AI tasks. In the AI engine, the task manager is the main component that coordinates and schedules the tasks and resources. It handles and parses the incoming AI tasks, and creates a dispatcher for each task. A dispatcher connects to multiple AI runtimes at external nodes. It also loads and caches the necessary data required by the corresponding AI tasks, performs data pipelines on it for preprocessing, feature engineering, etc, and pushes the prepared data and model weights to the remote AI runtime to trigger AI activities. Notably, the data is transferred in a streaming and pipelining manner to minimize the delay in the data preparation steps.
Data Streaming Protocol
During an AI task, the AI runtime receives data continuously from the database. NeurDB’s AI engine optimizes this process with a dedicated data streaming protocol to reduce the time and memory overheads. Specifically, the AI runtime establishes a TCP socket connection with the dispatcher. When a task is assigned to the dispatcher, it first schedules the AI runtimes and performs handshakes with them to negotiate (1) model parameters, such as model structure, model arguments, the training batch size, etc, and (2) streaming parameters, e.g., the initial size for send and receive buffers and the number of batches per transmission. Then, it starts the data and model transfer through the connection. Notably, to support adaptable control and scheduling over resources, these parameters can be dynamically updated for an ongoing AI task through a data-driven dispatcher. This makes the AI engine a self-driving component, thereby controlling and optimizing its operations autonomously.
Figure 3. Incremental Update for Model Manager
Model Manager
Given the dynamic nature of DBMSs, a typical AI lifecycle extends beyond a single model. As new data is introduced, incremental updates and retraining are required to address data and workload drifts that degrade predictive performance. This process results in multiple evolving model versions, creating significant management difficulties and storage overheads. To address this challenge, NeurDB leverages the capabilities of databases to efficiently manage AI models and handle drift by design. Specifically, it introduces a dedicated model manager in its AI engine, enabling fine-grained model management with efficient updates. The model manager provides high-level interfaces for handling AI operations, such as training, inference, and fine-tuning, executed via model views. Similar to data views in DBMSs, model views serve as logical abstractions of AI models tailored for specific tasks, with physical representations maintained in model storage. Formally, a model $M$ is uniquely defined by its Model ID (MID) $i$ and the timestamp of creation $t$, i.e., $M_{i,t}$. A model $M_{i,t}$, created by either users or internal components, comprises a series of layers $L$, each identified by a Layer ID (LID) $j$ and its training timestamp. To generate outputs for data $X$, the layers are executed sequentially: $M_{i,t}(X)=L^{(n)}(L^{(n-1)}(...(L^{(1)}(X)))$. This layered model storage approach aligns with the structure of deep neural networks (DNNs), ensuring fast and efficient AI model accessibility. Notably, this model formulation can also be generalized to non-linear DNN architectures that are represented as directed acyclic graphs, which can be achieved by executing layers based on the topological order.
Model Incremental Update
Leveraging its layered model storage, NeurDB enables linear versioning for a specific model and supports incremental model updates through fine-tuning. In particular, to adapt the current model to drifting data distributions, the AI engine selectively fine-tunes the final layers using the updated data stored in the designated database relation and freezes the preceding layers. Subsequently, only the updated layers remain persistent in the model storage, from where they can be extracted and merged with the previously frozen layers to create a new model version.
Figure 3 illustrates how the model manager supports incremental updates for AI models. Whenever an inference or fine-tuning task is initiated, the system retrieves all layers up to the most recent version to construct the model, i.e., for a model $M_{i,t}$ with $k$ layers, $M_{i,t}(X)=L_{i,t_{k}}^{(k)}(L_{i,t_{k-1}}^{(k-1)}(...(L_{i,t_{1}}^{(1)}(X))))\;s.t.\;\forall p,q\in\{1,...,k\},L_{i,t_{p}}^{(p)}\rightarrow t_{p}\geq t_{q}\wedge t_{p}\leq t.$ Let us consider the example shown in Figure 3. Supposing the model $M_{1}$ requires fine-tuning the last layer $L_{n}$, $M_{1,2}$, namely the second version of $M_{1}$, can then be assembled by layers $\{L_{1,1}^{(1)},...,L_{1,1}^{(n-1)},L_{1,2}^{(n)}\}$. This allows $M_{1,1}$ and $M_{1,2}$ to share the majority of model weights, ensuring adaptability to data drift while maintaining efficient storage.
4.2. Fast-adaptive Learned Components
Figure 4. Fast-adaptive Learned Concurrency Control
In this section, we introduce two key learned system components, including learned concurrency control and learned query optimizer, which can achieve fast adaptation to data and workload drift. To achieve this, we fundamentally rely on the underlying AI ecosystem. We non-intrusively monitor the system conditions such as transaction throughput and data distributions, which can be used to detect data and workload drift in real time. Based on these metrics, NeurDB automatically triggers model fine-tuning to adapt to continuously evolving data and workloads, and continually generates valid input for model pre-training, allowing the model for learned system components to gain global knowledge of most drifts.
Learned Concurrency Control
We design an efficient learned concurrency control algorithm that can continually adapt to changing workloads, as shown in Figure 4. Given a transaction $T$ with multiple operations, our algorithm incorporates a decision-making model $\mathcal{{F}}$ to assign each operation $op$ the optimal concurrency control action $\delta$ based on the current system condition represented as $x$, i.e., $\delta=\mathcal{{F}}(x)$. Unlike the state-of-the-art approach Wang et al. (2021) that simply adjusts actions based on predefined transaction or operation patterns (e.g., transaction type), our approach learns the optimal action based on the contention state, which includes both conflict information (such as dependency) of transactions and contextual information (such as the transaction length). For instance, when a write is performed on high-contention data records, we may immediately abort the transaction to avoid unnecessary costs, as the transaction is likely to be aborted eventually. In contrast, we may execute a read on low-contention records with optimistic concurrency control without locking to avoid extra conflict detection overhead, and a long transaction that is supposed to run for a long time should have high priority. By using a modelling paradigm based on the contention state, our model is more likely to generalize to drifting workloads with varying levels of contention. However, since transactions can be completed in milliseconds, the model must be efficient so as not to become a bottleneck. To achieve this, we first develop a fast encoding technique to significantly reduce the dimension of contention state representation, and then compress the model with a flattened layer to improve inference efficiency.
As the model is compressed to trade generalizability for performance, the representation of the contention state may gradually become inaccurate over time due to workload drift. Consequently, when we detect workload drift through their impact on transaction performance, we introduce a fine-tuning process for $\mathcal{{F}}$ to adapt it into $\mathcal{F}_{next}$. With the leaner architecture of the model, the adaptation can be accelerated due to the narrower search space compared to a large model. Specifically, we propose a two-phase adaptation algorithm based on the online Reinforcement Learning (RL) framework. In the first filtering phase, we generate several improved models using Bayesian optimization and evaluate them over a specific timeframe to identify the best-performing model. Then, in the refinement phase, we employ reward-based feedback to further optimize the selected model.
Figure 5. Fast-adaptive Learned Query Optimizer
Learned Query Optimizer
We propose a learned query optimizer that can efficiently adapt to data and workload drift, as shown in Figure 5. In contrast to existing works Marcus et al. (2021); Zhu et al. (2023) which attempt to produce the best plan for a given query $\mathcal{Q}$ under fixed system conditions (i.e., data distribution and workload), our approach effectively identifies the plan best suited for the current system conditions. To achieve this, we design a dual-module model consisting of an encoder and an analyzer. Specifically, in the encoder, we input the vector generated by a tree transformer, which includes multiple candidate query plans for $\mathcal{{Q}}$, along with system condition representations (including buffer information depicting buffer usage and data statistics representing each attribute’s distribution), into cross-attention layers to generate a unified embedding. The analyzer then uses this embedding as input for a multi-head attention layer followed by a multilayer perceptron (MLP) to deliver the optimal plan. With more knowledge of the mapping between plans and system conditions, the model can better generalize to data and workload drifts. To maximize this knowledge, we generate various synthetic data distributions and workloads using Bayesian optimization, and pre-train the model to handle most drifts effectively. Consequently, the proposed learned query optimizer provides consistent query performance under evolving data and workloads.
5. Evaluation
In this section, we present our preliminary evaluation results. We first introduce the experimental setup, and then evaluate the in-database AI ecosystem under real-world AI analytics scenarios, and the learned system components with data and workload drifts.
5.1. Experimental Setup
We conduct experiments on a server equipped with an Intel(R) Xeon(R) W-2133 CPU@3.60GHz (12 cores, 24 threads), 64 GB memory, and 3 NVIDIA GeForce RTX 2080 Ti GPUs. All experiments are executed within Docker containers based on the official Ubuntu 22.04 image with CUDA 11.8.0, leveraging the host’s GPU resources.
5.1.1. Benchmarks
We construct two real-world applications that require AI analytics.
E-commerce (E) Workload. We perform click-through rate prediction, a critical task in e-commerce for product recommendations, using the Avazu dataset (Avazu) ava (2024), which consists of ~40.4M records and 22 attributes. We use k-means clustering to create five data clusters namely, $C_{1}$ to $C_{5}$, and by switching from one to another, we simulate the data distribution drift.
Healthcare (H) Workload. We conduct disease progression prediction using the UCI Diabetes dataset (Diabetes) dia (2024). After scaling, the dataset comprises ~5.2M data records and 43 attributes.
We establish a micro-benchmark to evaluate the learned system components. It consists of a transactional benchmark based on YCSB Cooper et al. (2010), which generates synthetic workloads for large-scale Internet applications. Each transaction performs 5 selects and 5 updates on a table with 1 million records. In addition, we construct an OLAP benchmark based on the STATS dataset sta (2024), which consists of 8 tables from the Stats Stack Exchange network. We execute inserts/updates/deletes with randomly generated data values to simulate data distribution drift following a recent work Li et al. (2023b).
5.1.2. Implementation and Default Configuration
We have released the first version of NeurDB in neu (2024), in which we integrate our proposed in-database AI ecosystem and learned components into the codebase of PostgreSQL v16.3. We will gradually introduce new modules such as AI-powered resource scheduling, etc., and replace existing components where necessary. We implement a baseline system called PostgreSQL+P, which loads data from PostgreSQL in batches, and utilizes an AI runtime built with PyTorch to support AI analytics. By default, we employ ARM-Net Cai et al. (2021), an adaptive relation modelling network tailored for structured data, as the basic analytics model for both PostgreSQL+P and NeurDB. In our experiments, we inherit the default settings of PostgreSQL unless otherwise specified. We set the default window size of the streaming data loader to 80 data batches. Each batch contains 4096 data records (samples).
(a) Overall Performance
(b) Effects of Data Volume
(c) Effects of Data Distribution Drift
Figure 6. End-to-end Performance of AI Analytics
5.2. In-database AI Analytics
We evaluate the in-database AI analytics in terms of end-to-end latency, training throughput, and loss variation with data drift.
Efficiency and Scalability
We study the efficiency of NeurDB on AI analytics workloads by comparing it with PostgreSQL+P. As observed in Figure 6(a), NeurDB achieves up to 41.3% and 48.6% lower end-to-end latency, and 1.96$\times$ and 2.92$\times$ higher training throughput than PostgreSQL+P for Workload E and Workload H, respectively. The significant improvement in latency and throughput achieved by NeurDB is due to its in-database AI ecosystem, which efficiently supports AI analytics by utilizing a data streaming protocol. Further, to evaluate the impact of data volume on the end-to-end latency, we run Workload E with varying numbers of data batches. As shown in Figure 6(b), NeurDB consistently outperforms PostgreSQL+P, indicating that NeurDB can scale well with increased data volume.
Adaptability
We investigate NeurDB’s ability to adapt to the drifting data and workloads with the model incremental update technique. To simulate such drift, for $i\in[1,4]$, we let NeurDB perform the training task using cluster $C_{i}$ of Workload E, and switch to $C_{i+1}$ when 81,920 samples of $C_{i}$ are consumed by model training. Figure 6(c) plots the training losses with and without model incremental updates. From the result, we can observe that starting from the first data drift, the AI engine equipped with incremental updates receives lower loss values during the sudden drift in data distributions. This enables the model to converge faster, and as a result, NeurDB is equipped to serve the new tasks effectively.
5.3. Learned System Components
We now investigate the performance of our proposed learned system components.
(a) Overall Throughput      (b) Throughput under Data and Workload Drift
Figure 7. Performance of Learned Concurrency Control
Learned Concurrency Control
To evaluate the proposed concurrency control algorithm, we compare NeurDB with PostgreSQL using the micro-benchmark with varying thread counts. The results, shown in Figure 7(a), demonstrate that NeurDB achieves up to 1.44$\times$ higher transaction throughput than PostgreSQL. We attribute this performance gain to the proposed algorithm’s ability to schedule transactions more effectively than serializable snapshot-isolation Ports and Grittner (2012), a static concurrency control algorithm employed in PostgreSQL. We further evaluate the adaptability of our proposed algorithm against Polyjuice Wang et al. (2021), the state-of-the-art learned algorithm. Due to the limitation of constraining the transaction execution workflow, implementing Polyjuice on the codebase of NeurDB would be cumbersome. We therefore opt to implement our algorithm, named NeurDB(CC), into the Polyjuice codebase to facilitate a fair comparison. We set up a drift workload based on TPCC Council (2007) by varying the number of warehouses and threads. As shown in Figure 7(b), NeurDB(CC) adapts quickly to workload drifts and outperforms Polyjuice by up to 2.05$\times$. The superior performance of NeurDB(CC) mainly stems from its design that encapsulates a fast yet accurate model to find the best concurrency control action, while facilitating the fine-tuning process with the two-phase adaptation algorithm.
Learned Query Optimizer
We next compare NeurDB with PostgreSQL and two state-of-the-art learned query optimizer approaches, namely Bao Marcus et al. (2021), and Lero Zhu et al. (2023). We construct three workloads with different data distributions and randomly select 8 SPJ queries provided by STATS datasets. We use stable models of Bao and Lero for the experiment, as they demonstrated good performance in their respective papers.
As can be observed in Figure 8, NeurDB achieves up to 20.32% lower average latency of all evaluated queries, which demonstrates its effective adaptability to both data and workload drifts. Due to the proposed dual-module model, NeurDB is able to effectively capture system conditions and use them to select an efficient query plan.
Figure 8. Performance of Learned Query Optimizers
6. Related Work
Our work relates to a broad spectrum of efforts on the fusion of AI and databases. The initial concepts Brodie (1988); Graefe and DeWitt (1987) can be traced back to the 1980s. At that time, the fusion was far from reality due to limited development in both realms. With advancements in both fields over the years, numerous attempts by academia  Kraska et al. (2019); Pavlo et al. (2017) and industry Karanasos et al. (2020); Li et al. (2021) have been made. However, many problems remain open, and it is important to consider database dynamicity in the fusion of AI with databases. Further, the system architecture may require a re-design to support AI-powered operations. NeurDB has been designed along this direction by seamlessly integrating AI into the database architecture, enabling continuous adaptation to data and workload drift.
7. Conclusions
This paper presents NeurDB, a novel AI-powered autonomous DBMS that is adaptable to data and workload drift. NeurDB, with its fast-adaptive learned system components and in-database AI ecosystem, facilitates efficient and effective in-database AI analytics. Empirical evaluations demonstrate the superiority of NeurDB, highlighting its potential to realize a seamless fusion of AI and databases.

### v2

@@
 Abstract.
-Databases are increasingly embracing AI to provide autonomous system optimization and intelligent in-database analytics, aiming to relieve end-user burdens across various industry sectors.
+Databases are increasingly embracing AI to provide intelligent in-database analytics and autonomous system optimization, aiming to relieve end-user burdens across various industry sectors.
 Nonetheless, most existing approaches fail to account for the dynamic nature of databases, which renders them ineffective for real-world applications characterized by evolving data and workloads.
@@
 NeurDB establishes a new in-database AI ecosystem that seamlessly integrates AI workflows within the database.
-This integration enables efficient and effective in-database AI analytics and fast-adaptive learned system components.
-Empirical evaluations demonstrate that NeurDB substantially outperforms existing solutions in managing AI analytics tasks, with the proposed learned components more effectively handling environmental dynamism than state-of-the-art approaches.
+This integration enables efficient and effective in-database AI analytics and fast-adaptive learned database components.
+Empirical evaluations demonstrate that NeurDB substantially outperforms existing solutions in managing AI analytics tasks, with the proposed learned components more effectively handling data and workload drift than state-of-the-art approaches.
 1. Introduction
-Database management systems (DBMSs) are becoming more automated and intelligent by embracing artificial intelligence (AI).
-AI empowers these systems to achieve autonomous optimization Kraska et al.
-(2019); Pavlo et al.
-(2017); Zhu et al.
-(2024) and support complex analytics Karanasos et al.
+Database management systems (DBMSs) are becoming more intelligent and autonomous by embracing artificial intelligence (AI), catering to modern data- and AI-centric applications.
+AI augments DBMS functionality to support in-database AI analytics Karanasos et al.
 (2020); Xing et al.
 (2024); Ooi et al.
-(2015) with minimal human intervention, thereby catering to advanced data- and AI-centric applications.
+(2015); Wang et al.
+(2015), enabling complex and advanced analytics tasks such as disease progression predictions and user purchase recommendations.
+In parallel, AI empowers DBMSs to achieve autonomous DBMS optimization, driving a broad range of learned database designs that enhance system performance with minimal human intervention, including learned knob tuners Pavlo et al.
+(2017), learned system drivers Zhu et al.
+(2024); Li et al.
+(2021), and learned database components Kraska et al.
+(2019) such as query optimizers Marcus et al.
+(2021); Zhu et al.
+(2023), indexes Kraska et al.
+(2018), and concurrency control Wang et al.
+(2021), etc.
 The aspiration of integrating DBMSs with AI was first expressed forty years ago Brodie (1988), which has been periodically revisited with evolving technology Wang et al.
@@
 With the advancements in both AI and DBMSs, considerable progress has been made in deepening their fusion.
-Unfortunately, a huge gap remains between the potential of this integration and its current state of usability, largely due to the inherent differences in their paradigms.
-The dynamic nature of databases, characterized by constantly changing data and workloads due to transactional updates, and user interactions, poses a fundamental challenge.
-For example, an e-commerce database might experience a sharp increase in workloads during flash sales compared to normal periods, caused by huge volumes of updates from sales transactions and processing.
+Unfortunately, a huge gap remains between the potential of this integration and its current state of usability, largely due to the inherent differences in their paradigms Ooi et al.
+(2024).
+The dynamic nature of databases, characterized by data and workload drift, poses a fundamental challenge Wu and Ives (2024); Kurmanji et al.
+(2024); Kurmanji and Triantafillou (2023).
+For example, suppose an e-commerce database stores user and product information.
+While the database continuously evolves due to user and merchant activities in normal periods, it would experience a sharp increase in workloads during flash sales, where huge volumes of updates from sales transactions are processed.
 In contrast, AI models typically derive intelligence from static datasets and thus can become outdated quickly in the face of database dynamism.
-For example, models for learned query optimization often acquire knowledge by training on specific system environments including workload patterns and data distributions.
-However, when these environments drift, e.g., due to data updates from user interactions, models relying on past knowledge may lead to ineffective query optimizations since they are not updated to reflect new conditions.
-The discrepancy in dynamicity between AI and DBMSs poses a great challenge to the effective adoption of AI-enhanced DBMSs.
-First, traditional AI-driven system optimization, including knob tuning Pavlo et al.
-(2017), intelligent advisors Zhu et al.
-(2024); Li et al.
-(2021), and learned system components Kraska et al.
-(2019) such as query optimizers Marcus et al.
-(2021); Zhu et al.
-(2023), indexes Kraska et al.
-(2018), and concurrency control Wang et al.
-(2021), often overlook the dynamicity of data distributions and workloads.
-Second, another important facet of AI and DBMS integration, namely in-database AI analytics for complex tasks such as disease progression predictions and user purchase recommendations, has gained traction but faces similar adaptability issues.
-Existing systems generally focus on static model inference Salazar-Díaz et al.
-(2024); Park et al.
-(2022), model serving Vartak (2017), and others Xu et al.
-(2022), without considering necessary model adaptations to evolving data and workload patterns.
-Consequently, they require frequent manual updates and retraining to remain effective, which complicates the analytics workflow.
-Recent works Wu and Ives (2024); Li et al.
-(2022, 2023a); Kurmanji and Triantafillou (2023) aim to address these challenges but are typically more effective in controlled scenarios, such as data distribution changes following certain rules or query optimization under read-only workloads.
-These methods have not been designed to generalize across more dynamic and practical workloads involving continuous read and write operations.
-We envision a deep integration of AI and DBMSs for adaptability to both data and workload drift.
+For instance, models for user purchase recommendations typically acquire knowledge by training on specific user behaviors and product information.
+However, when the data drifts due to transactional updates, models relying on past knowledge may produce inaccurate recommendations since they are not updated to reflect new conditions.
+Similarly, learned query optimizers trained based on historical system environments including data distributions and workload patterns can struggle to deliver effective query plans as continuous data and workload drift.
+Adapting to data and workload drift has gained traction for both in-database AI analytics and autonomous DBMS optimization.
+Existing databases supporting in-database AI analytics SQL (2024); PostgresML (2024); MindsDB (2024) often recommend users to completely retrain models with new data and workloads when the models become outdated.
+This retraining process is typically performed manually outside the database, and therefore, it may complicate the AI analytics workflow and becomes inefficient when dealing with continuous data and workload drift.
+To tackle this problem, in the context of autonomous DBMS optimization, recent works Wu and Ives (2024); Li et al.
+(2022, 2023a); Kurmanji and Triantafillou (2023) have started addressing the adaptability issues by automatically detecting drift and triggering model retraining.
+However, these approaches are often tailored to certain components.
+For example, approaches Wu and Ives (2024); Kurmanji and Triantafillou (2023) that allow learned query optimizers to withstand workload pattern changes can be ineffective for learned indexes requiring rapid adaptation to data distribution drift.
+Consequently, a paradigm that robustly handles data and workload drift for various learned database components remains absent.
+We envision a deeper integration of AI and DBMSs for enhanced adaptability to both data and workload drift.
 Unlike existing approaches that merely overlay AI onto DBMSs or selectively enhance certain system components, we aim for a comprehensive fusion of AI with DBMSs that enables continuous adaptation to drift.
+It promises a cohesive AI-powered DBMS that provides advanced in-database AI analytics and deeply integrates AI into all its key components.
 However, achieving this level of integration presents significant challenges.
 First, the integration necessitates a fundamental redesign of the entire AI workflow, including training, inference, and fine-tuning, within the database architecture.
-Second, models for learned database components and AI analytics must swiftly adapt to drifting data workloads without losing effectiveness and efficiency.
-Therefore, developing such a system requires both a system foundation that seamlessly supports model adaptation and AI models that are inherently adaptive in structure.
-In this paper, we present NeurDB [^2] https://github.com/neurdb/neurdb, an AI-powered autonomous DBMS that provides efficient and effective in-database AI analytics to seamlessly support modern AI-powered applications, and fully embraces AI techniques in each major system component to offer self-driving capabilities.
-At the core of NeurDB is an in-database AI ecosystem that achieves deep integration of the AI workflow into the database.
-In this ecosystem, we develop multiple in-database AI operations, such as model training, inference, and fine-tuning, along with an in-database AI engine to handle the execution of these operators.
-NeurDB can then directly support in-database AI analytics by calling these AI operations.
-To further simplify end-users in submitting their complex AI analytics tasks to NeurDB, we provide a user-friendly interface by extending SQL syntax with PREDICT syntax.
+Second, models for AI analytics and learned database components must swiftly adapt to data and workload drift without losing effectiveness and efficiency.
+Therefore, developing such an AI-powered database requires both a system foundation that seamlessly supports model adaptation and AI models that are inherently adaptive in structure.
+In this paper, we present NeurDB, an AI-powered autonomous DBMS that provides efficient and effective in-database AI analytics to seamlessly support modern AI-powered applications, and fully embraces AI techniques in each major system component to offer autonomous system optimization.
+At the core of NeurDB is an in-database AI ecosystem that deeply integrate the AI workflow into the database.
+In this ecosystem, we develop multiple in-database AI operators, such as model training, inference, and fine-tuning, along with an in-database AI engine to handle the execution of these operators.
+NeurDB can then directly support in-database AI analytics by calling these AI operators.
+To further simplify end-users in submitting their AI analytics tasks to NeurDB, we provide a user-friendly interface by extending SQL with PREDICT syntax.
 Moreover, we propose two techniques to optimize the AI ecosystem performance.
@@
 Second, we develop an incremental update technique to minimize the fine-tuning cost, facilitating the fast adaptation of AI models.
-Based on this foundation, we propose two efficient learned system components, namely a learned concurrency control algorithm and a learned query optimizer, that can adapt quickly to data and workload drift.
-2. Preliminaries
-In this section, we outline the design goals of NeurDB and present the SQL syntax enhanced for AI-powered Analytics.
+Based on the in-database AI ecosystem, we enable two efficient learned database components, namely a learned concurrency control algorithm and a learned query optimizer, that can adapt quickly to data and workload drift.
+The remainder of the paper is structured as follows.
+The next section provides the necessary preliminaries and further clarifies the motivation.
+Section 3 describes the system overview of NeurDB, and Section 4 details the key system design, including the in-database AI ecosystem and fast-adaptive learned database components.
+Section 5 presents the experimental results.
+Section 6 discusses the related works before Section 7 concludes.
+2. Preliminaries and Motivation
+In this section, we introduce the preliminaries of AI and DBMS integration, outline the design goals of NeurDB, and present the enhanced SQL syntax designed for in-database AI analytics.
+2.1. AI and DBMS Integration
+The integration of AI and DBMSs can offer each other mutual benefits Wang et al.
+(2016), and building on this intuition, various research works on AI and DBMS integration have been proposed, which can be broadly categorized into two areas: in-database analytics and autonomous system optimization.
+Modern databases, such as Oracle Oracle (2024), Microsoft Azure SQL SQL (2024), Amazon Redshift Redshift (2024), and Google BigQuery BigQuery (2024), support in-database AI analytics, and allow AI tasks to be performed directly through user-friendly SQL interfaces.
+For better performance, these databases enable AI-driven knob tuning to autonomously refine database configurations.
+Adapting to data and workload drift has been recognized as a critical challenge in advancing the integration of AI and DBMSs Wu and Ives (2024); Kurmanji et al.
+(2024); Kurmanji and Triantafillou (2023).
+Many existing approaches for whether in-database analytics Karanasos et al.
+(2020); Salazar-Díaz et al.
+(2024) or autonomous system optimization Marcus et al.
+(2019); Kraska et al.
+(2018) depend on static models or complete retraining, limiting their ability to handle continuous data and workload drift.
+More importantly, these approaches are typically layered on top of DBMSs rather than being holistically embedded within DBMSs.
+This separation hinders real-time adaptability, because internal and fine-grained performance metrics, which are essential for detecting and addressing drift, can be difficult to capture externally Wang et al.
+(2021); Li et al.
+(2022).
+To mitigate these limitations, we design and develop NeurDB, an AI-powered database with AI holistically embedded from the ground up.
+By deeply integrating AI into all its key components and functionalities, NeurDB aims to automatically manage adaptation workflows for both in-database AI analytics and autonomous system optimization.
 2.2. Design Goals
-DBMSs are dynamic as data distributions and workloads evolve over time, and therefore, the system must be designed for adaptability while also guaranteeing reliability and scalability.
-Building upon this understanding, we define three essential properties for NeurDB.
+DBMSs are dynamic as data and workloads evolve over time, and therefore, the system must be designed for adaptability while also guaranteeing reliability and scalability.
+Building upon this understanding, we equip NeurDB with these three key properties.
 Adaptability is the capability of a DBMS to evolve autonomously in response to drifting data and workloads. With optimal adaptability, DBMS can respond to drift in real time.
@@
 Scalability refers to the system’s capacity to maintain or enhance performance as the workload increases by introducing more resources, such as threads or nodes.
-Informally, the design goal of NeurDB is to ensure the system can uphold reliability as quickly as possible when the system or its components adapt or evolve due to data and workload drift, while ensuring good scalability.
+Informally, our design goal is to ensure NeurDB can uphold reliability as quickly as possible when the system adapts or evolves due to data and workload drift, while ensuring good scalability.
+In addition, we design NeurDB as a general-purpose AI-powered DBMS which is expected to serve as a foundational infrastructure for advanced data- and AI-centric applications to enrich various domains with AI capabilities Ooi et al.
+(2024).
 2.3. SQL Syntax for AI Analytics
-NeurDB incorporates enhanced SQL, which extends from the standard SQL to support AI analytics.
-As illustrated in Listing 1 and Listing 2, NeurDB introduces a new PREDICT keyword to handle two typical ML tasks: classification with the CLASS OF clause and regression with the VALUE OF clause.
-Inspired by the original principle of introducing SQL that allows an application developer to write a query with SELECT, and then let the DBMS find the most efficient way to store and retrieve data, we ensure that a developer can submit an AI analytic task simply with PREDICT.
-All the following operations, such as retrieving training data and invoking AI models, are handled automatically by the database.
-We also plan to expose more enhanced SQL for AI model management and other AI analytics functionalities.
-Due to space constraints, we mainly focus on the PREDICT query in this paper.
-We now present two real-world analytic scenarios that can be directly supported by NeurDB using PREDICT.
+NeurDB incorporates enhanced SQL to support AI analytics.
+As illustrated in Listing 1 and Listing 2, it extends from the standard SQL by introducing a PREDICT keyword to handle two typical AI tasks: regression with the VALUE OF clause and classification with the CLASS OF clause.
+Inspired by the original principle of SQL that allows an application developer to write a query with SELECT, and then let the DBMS find the most efficient way to store and retrieve data, we ensure that a developer can submit an AI analytics task simply with PREDICT.
+All the following operations, such as retrieving training data and invoking AI models, are handled automatically by NeurDB.
+Consequently, unlike existing solutions PostgresML (2024); BigQuery (2024) that require application developers to intervene in the execution of AI analytics, such as specifying model parameters, our approach can make the execution transparent to application developers, relieving them from the complexities of performing such analytics.
+Further, we plan to expose additional SQL for AI model management and other AI analytics functionalities, such as generative AI.
+We now present two real-world analytics scenarios that can be directly supported by NeurDB using PREDICT queries.
+Regression
+Listing 1 shows the SQL to predict and fill in the missing scores of products based on user reviews.
+By specifying the features for training with TRAIN ON and using data stored in the ‘review’ table, we predict the value of the ‘score’ variable.
+Notably, for TRAIN ON, the asterisk automatically excludes features with unique constraints to avoid including meaningless data.
+Listing 1: Syntax for a regression task
+1PREDICT VALUE OF score
+2FROM review
+3WHERE brand_name = ’SpecialGoods’
+4TRAIN ON *
+5 WITH brand_name <> ’SpecialGoods’
 Classification
-The disease progression prediction in healthcare can be handed by the SQL shown in Listing 1.
-By specifying the features with TRAIN ON and using data stored in the ‘diabetes’ table and directly input with VALUES, we predict the class of the ‘outcome’ variable, which indicates whether a patient has diabetes or not.
-Regression
-Listing 2 shows the SQL to predict and fill in the missing scores of products based on user reviews.
-Similarly, it specifies the features for training and the data table to predict the value of the ‘score’ variable.
-Notably, it automatically excludes features with unique constraints to avoid including meaningless data.
-Listing 1: Syntax for a classification task
+The disease progression prediction in healthcare can be handed by the SQL shown in Listing 2.
+Similarly, this query specifies the features for training with TRAIN ON and the data table using FROM to predict the class of the ‘outcome’ variable, which indicates whether a patient has diabetes or not.
+In addition, we support directly inputting missing data using VALUES.
+Listing 2: Syntax for a classification task
 1PREDICT CLASS OF outcome
@@
 4VALUES (6, 148, 72, ...), (1, 85, 66, ...), ...
-Listing 2: Syntax for a regression task
-1PREDICT VALUE OF score
-2FROM review
-3WHERE brand_name = ’SpecialGoods’
-4TRAIN ON * WITH brand_name <> ’SpecialGoods’
 3. System Overview
-Figure 1. The System Architecture of NeurDB
 We now describe the system architecture of NeurDB as shown in Figure 1, where we achieve the deep fusion of AI and databases by establishing an in-database AI ecosystem.
-To achieve this, we first holistically redesign each existing database component, such as the query optimizer, transaction manager, and index manager, to support in-database AI analytics and function as learned components for improved performance.
-In particular, we enhance the SQL parser and the query optimizer to produce a customized query plan for PREDICT queries.
+Based on this AI ecosystem, we enable efficient and effective in-database AI analytics and develop fast-adaptive learned database components.
+In-database AI Ecosystem
+We establish the in-database AI ecosystem by holistically redesigning existing database components, such as the query executor, and introducing additional modules, including an AI engine and AI model storage.
+We first extend the query executor to support AI workflow.
 Beyond traditional operators for fetching and processing data, e.g., scan and join, the query executor in NeurDB includes in-database AI operators for model training, inference, and fine-tuning.
-With these AI operators, we effectively align the AI workflow with the database query processing.
-We are developing additional AI operators, such as model selection, to provide more comprehensive AI services.
-For example, a query may call the model selection operator (denoted as MSelection) to automatically select the best-suited model for a given prediction task Xing et al.
-(2024), thereby enhancing accuracy and efficiency.
-We then introduce a new component, called the AI engine, into NeurDB.
+With these AI operators, we effectively integrate the AI workflow into the database query processing.
+We are developing additional AI operators, such as model selection Xing et al.
+(2024) and model slicing Zeng et al.
+(2025), to provide more comprehensive AI services.
+For example, a query may call the model selection operator (denoted as MSelection) to automatically select the best-suited model for a given prediction task, thereby enhancing accuracy and efficiency.
+We then introduce a new module, called the AI engine, into NeurDB.
 It handles the AI-related processing requests from AI operators and learned components and creates AI tasks on their behalf, where the task manager dispatches them to the CPU/GPU runtime for execution.
@@
 We further implement a monitor to detect unexpected performance or accuracy issues, based on which we trigger automatic and appropriate model adaptation.
+In-database AI Analytics
 In Figure 1, we illustrate a running example of a PREDICT query.
-After parsing and optimizing the query, the executor performs the scan operation to retrieve data and then invokes the inference operator to deliver the inference task to the AI engine.
+We enhance the SQL parser and the query optimizer to produce a customized query plan for PREDICT queries.
+After parsing and optimizing the query, the query executor then executes the query according to the query plan.
+In particular, it performs the scan operator to retrieve data and then invokes the inference operator to deliver the inference task to the AI engine.
 Next, the AI engine uploads the model from the model storage if it is not in the model buffer, and conducts model inference to produce the results.
+If the model is detected to be inaccurate, NeurDB invokes the fine-tuning operator to update the outdated model with the help of the AI engine.
+Autonomous DBMS Optimization
+We also enable autonomous system optimization based on the proposed in-database AI ecosystem.
+For instance, the learned query optimizer interacts with the AI engine to train a model responsible for generating efficient query plans.
+The monitor tracks the performance of the generated query plans, and if a plan is identified as inefficient, the monitor notifies the AI engine to fine-tune the model, enabling continuous optimization and adaptability.
+Similarly, the monitor can trigger autonomous knob tuning when suboptimal knob settings are detected, ensuring that the system remains well-configured to handle data and workload drift effectively.
+We will further introduce the detailed design of learned database components in Section 4.2.
+Figure 1. The System Architecture of NeurDB
 4. The Design of NeurDB
-In this section, we present two key subsystems of NeurDB, namely an in-database AI ecosystem and learned system components.
+In this section, we present two key subsystems of NeurDB, namely an in-database AI ecosystem and learned database components.
 4.1. In-database AI Ecosystem
-Figure 2. AI Engine of NeurDB
 AI Engine
@@
 Notably, the data is transferred in a streaming and pipelining manner to minimize the delay in the data preparation steps.
+Figure 2. AI Engine of NeurDB
 Data Streaming Protocol
@@
 This makes the AI engine a self-driving component, thereby controlling and optimizing its operations autonomously.
-Figure 3. Incremental Update for Model Manager
 Model Manager
 Given the dynamic nature of DBMSs, a typical AI lifecycle extends beyond a single model.
-As new data is introduced, incremental updates and retraining are required to address data and workload drifts that degrade predictive performance.
+As new data is introduced, incremental updates and retraining are required to address data and workload drift that degrade predictive performance.
 This process results in multiple evolving model versions, creating significant management difficulties and storage overheads.
-To address this challenge, NeurDB leverages the capabilities of databases to efficiently manage AI models and handle drift by design.
+To address this challenge, NeurDB leverages the capabilities of databases to efficiently manage AI models created by either users or internal components and handle drift by design.
 Specifically, it introduces a dedicated model manager in its AI engine, enabling fine-grained model management with efficient updates.
@@
 Similar to data views in DBMSs, model views serve as logical abstractions of AI models tailored for specific tasks, with physical representations maintained in model storage.
-Formally, a model $M$ is uniquely defined by its Model ID (MID) $i$ and the timestamp of creation $t$, i.e., $M_{i,t}$.
-A model $M_{i,t}$, created by either users or internal components, comprises a series of layers $L$, each identified by a Layer ID (LID) $j$ and its training timestamp.
-To generate outputs for data $X$, the layers are executed sequentially: $M_{i,t}(X)=L^{(n)}(L^{(n-1)}(...(L^{(1)}(X)))$.
+Formally, a model $M$ comprises a series of layers $L^{(j)}(j\in\mathbb{Z}^{+})$.
+To generate outputs for data $X$, the layers are executed sequentially, i.e., $M(X)=L^{(n)}(L^{(n-1)}(...(L^{(1)}(X)))$.
 This layered model storage approach aligns with the structure of deep neural networks (DNNs), ensuring fast and efficient AI model accessibility.
 Notably, this model formulation can also be generalized to non-linear DNN architectures that are represented as directed acyclic graphs, which can be achieved by executing layers based on the topological order.
+Figure 3. Incremental Update for Model Manager
 Model Incremental Update
-Leveraging its layered model storage, NeurDB enables linear versioning for a specific model and supports incremental model updates through fine-tuning.
+Leveraging its layered model storage, NeurDB enables versioning for a specific model and supports incremental model updates through fine-tuning.
 In particular, to adapt the current model to drifting data distributions, the AI engine selectively fine-tunes the final layers using the updated data stored in the designated database relation and freezes the preceding layers.
@@
 Figure 3 illustrates how the model manager supports incremental updates for AI models.
-Whenever an inference or fine-tuning task is initiated, the system retrieves all layers up to the most recent version to construct the model, i.e., for a model $M_{i,t}$ with $k$ layers, $M_{i,t}(X)=L_{i,t_{k}}^{(k)}(L_{i,t_{k-1}}^{(k-1)}(...(L_{i,t_{1}}^{(1)}(X))))\;s.t.\;\forall p,q\in\{1,...,k\},L_{i,t_{p}}^{(p)}\rightarrow t_{p}\geq t_{q}\wedge t_{p}\leq t.$ Let us consider the example shown in Figure 3.
+Whenever an inference or fine-tuning task is initiated, the system retrieves all layers up to the most recent version to construct the model, i.e., for a model $M_{i,t}$, uniquely defined by its Model ID (MID) $i$ and the timestamp of creation $t$, consisting of $k$ layers,
+$M_{i,t}(X)=L_{i,t_{k}}^{(k)}(L_{i,t_{k-1}}^{(k-1)}(...(L_{i,t_{1}}^{(1)}(X))))$
+$\;s.t.\;\forall p,q\in\{1,...,k\},L_{i,t_{p}}^{(p)}\rightarrow t_{p}\geq t_{q}\wedge t_{p}\leq t.$
+Let us consider the example shown in Figure 3.
 Supposing the model $M_{1}$ requires fine-tuning the last layer $L_{n}$, $M_{1,2}$, namely the second version of $M_{1}$, can then be assembled by layers $\{L_{1,1}^{(1)},...,L_{1,1}^{(n-1)},L_{1,2}^{(n)}\}$.
@@
 4.2. Fast-adaptive Learned Components
-Figure 4. Fast-adaptive Learned Concurrency Control
-In this section, we introduce two key learned system components, including learned concurrency control and learned query optimizer, which can achieve fast adaptation to data and workload drift.
+In this section, we introduce two key learned database components, learned concurrency control and learned query optimizer, which can achieve fast adaptation to data and workload drift.
 To achieve this, we fundamentally rely on the underlying AI ecosystem.
 We non-intrusively monitor the system conditions such as transaction throughput and data distributions, which can be used to detect data and workload drift in real time.
-Based on these metrics, NeurDB automatically triggers model fine-tuning to adapt to continuously evolving data and workloads, and continually generates valid input for model pre-training, allowing the model for learned system components to gain global knowledge of most drifts.
+Based on these metrics, NeurDB automatically triggers model fine-tuning to adapt to continuously evolving data and workloads, and continually generates valid input for model pre-training, allowing the model for learned database components to gain global knowledge of most drift.
 Learned Concurrency Control
@@
 Given a transaction $T$ with multiple operations, our algorithm incorporates a decision-making model $\mathcal{{F}}$ to assign each operation $op$ the optimal concurrency control action $\delta$ based on the current system condition represented as $x$, i.e., $\delta=\mathcal{{F}}(x)$.
-Unlike the state-of-the-art approach Wang et al.
+Unlike state-of-the-art approach Wang et al.
 (2021) that simply adjusts actions based on predefined transaction or operation patterns (e.g., transaction type), our approach learns the optimal action based on the contention state, which includes both conflict information (such as dependency) of transactions and contextual information (such as the transaction length).
@@
 In contrast, we may execute a read on low-contention records with optimistic concurrency control without locking to avoid extra conflict detection overhead, and a long transaction that is supposed to run for a long time should have high priority.
-By using a modelling paradigm based on the contention state, our model is more likely to generalize to drifting workloads with varying levels of contention.
+By using a modeling paradigm based on the contention state, our model is more likely to generalize to drifting workloads with varying levels of contention.
 However, since transactions can be completed in milliseconds, the model must be efficient so as not to become a bottleneck.
 To achieve this, we first develop a fast encoding technique to significantly reduce the dimension of contention state representation, and then compress the model with a flattened layer to improve inference efficiency.
+Figure 4. Fast-adaptive Learned Concurrency Control
 As the model is compressed to trade generalizability for performance, the representation of the contention state may gradually become inaccurate over time due to workload drift.
@@
 Then, in the refinement phase, we employ reward-based feedback to further optimize the selected model.
-Figure 5. Fast-adaptive Learned Query Optimizer
 Learned Query Optimizer
@@
 (2021); Zhu et al.
-(2023) which attempt to produce the best plan for a given query $\mathcal{Q}$ under fixed system conditions (i.e., data distribution and workload), our approach effectively identifies the plan best suited for the current system conditions.
+(2023) that attempt to produce the best plan for a given query $\mathcal{Q}$ under fixed system conditions (i.e., data distribution and workload), our approach effectively identifies the plan best suited for the current system conditions.
 To achieve this, we design a dual-module model consisting of an encoder and an analyzer.
@@
 The analyzer then uses this embedding as input for a multi-head attention layer followed by a multilayer perceptron (MLP) to deliver the optimal plan.
-With more knowledge of the mapping between plans and system conditions, the model can better generalize to data and workload drifts.
-To maximize this knowledge, we generate various synthetic data distributions and workloads using Bayesian optimization, and pre-train the model to handle most drifts effectively.
+With more knowledge of the mapping between plans and system conditions, the model can better generalize to data and workload drift.
+To maximize this knowledge, we generate various synthetic data distributions and workloads using Bayesian optimization, and pre-train the model to handle most drift effectively.
 Consequently, the proposed learned query optimizer provides consistent query performance under evolving data and workloads.
+Figure 5. Fast-adaptive Learned Query Optimizer
+Discussion
+The main design principle we follow to develop computational models and algorithms in NeurDB is the filter-and-refine principle (FRP) Ooi et al.
+(2024, 2002); Filter and Refine (2024).
+FRP employs a two-stage strategy: a filtering stage quickly eliminates less promising or irrelevant objects (e.g., data, strategies, etc.) from a large set, followed by a more resource-intensive refinement stage on the filtered subset.
+This strategy enables the efficient execution of computationally demanding tasks while yielding near-optimal results.
+We therefore leverage FRP in both in-database AI analytics and autonomous DBMS optimization for enhanced adaptability.
+For example, we integrate FRP in the RL framework of learned concurrency control and incorporate it in the multi-head attention layers of the learned query optimizer.
+However, existing model structures, typically characterized by fixed topologies and layer configurations, constrain the full potential of FRP.
+We are exploring a flexible model structure that can be dynamically adjusted with the guidance of FRP, enabling seamless adaptation to data and workload drift.
 5. Evaluation
 In this section, we present our preliminary evaluation results.
-We first introduce the experimental setup, and then evaluate the in-database AI ecosystem under real-world AI analytics scenarios, and the learned system components with data and workload drifts.
+We first introduce the experimental setup, and then evaluate the in-database AI ecosystem under real-world AI analytics scenarios, and the learned database components with data and workload drifts.
 5.1. Experimental Setup
@@
 5.1.1. Benchmarks
-We construct two real-world applications that require AI analytics.
-E-commerce (E) Workload.
-We perform click-through rate prediction, a critical task in e-commerce for product recommendations, using the Avazu dataset (Avazu) ava (2024), which consists of ~40.4M records and 22 attributes.
+We construct two real-world applications that require AI analytics. The AI analytics queries used in our experiments are listed in Table 1.
+  • E-commerce (E) Workload performs click-through rate prediction, a critical task in e-commerce for product recommendations, using the Avazu dataset (Avazu) Dataset (2024a), which consists of ~40.4M records and 22 attributes.
 We use k-means clustering to create five data clusters namely, $C_{1}$ to $C_{5}$, and by switching from one to another, we simulate the data distribution drift.
-Healthcare (H) Workload. We conduct disease progression prediction using the UCI Diabetes dataset (Diabetes) dia (2024). After scaling, the dataset comprises ~5.2M data records and 43 attributes.
-We establish a micro-benchmark to evaluate the learned system components.
+  • Healthcare (H) Workload conducts disease progression prediction using the UCI Diabetes dataset (Diabetes) Dataset (2024c).
+After scaling, the dataset comprises ~5.2M data records and 43 attributes.
+We establish a micro-benchmark to evaluate the learned database components.
 It consists of a transactional benchmark based on YCSB Cooper et al.
@@
 Each transaction performs 5 selects and 5 updates on a table with 1 million records.
-In addition, we construct an OLAP benchmark based on the STATS dataset sta (2024), which consists of 8 tables from the Stats Stack Exchange network.
+In addition, we construct an OLAP benchmark based on the STATS dataset Dataset (2024b), which consists of 8 tables from the Stats Stack Exchange network.
 We execute inserts/updates/deletes with randomly generated data values to simulate data distribution drift following a recent work Li et al.
-(2023b).
+(2023a).
+Table 1. Queries for AI Analytics Evaluations
+Workload Statement
+E-Commerce (E) PREDICT VALUE OF click_rate FROM avazu TRAIN ON *
+Healthcare (H) PREDICT CLASS OF outcome FROM diabetes TRAIN ON *
 5.1.2. Implementation and Default Configuration
-We have released the first version of NeurDB in neu (2024), in which we integrate our proposed in-database AI ecosystem and learned components into the codebase of PostgreSQL v16.3.
+We have released the first version of NeurDB NeurDB (2024), in which we integrate our proposed in-database AI ecosystem and learned components into the codebase of PostgreSQL v16.3.
 We will gradually introduce new modules such as AI-powered resource scheduling, etc., and replace existing components where necessary.
@@
 By default, we employ ARM-Net Cai et al.
-(2021), an adaptive relation modelling network tailored for structured data, as the basic analytics model for both PostgreSQL+P and NeurDB.
+(2021), an adaptive relation modeling network tailored for structured data, as the basic analytics model for both PostgreSQL+P and NeurDB.
 In our experiments, we inherit the default settings of PostgreSQL unless otherwise specified.
@@
 This enables the model to converge faster, and as a result, NeurDB is equipped to serve the new tasks effectively.
-5.3. Learned System Components
-We now investigate the performance of our proposed learned system components.
+5.3. Learned Database Components
+We now investigate the performance of our proposed learned database components.
 (a) Overall Throughput      (b) Throughput under Data and Workload Drift
@@
 The results, shown in Figure 7(a), demonstrate that NeurDB achieves up to 1.44$\times$ higher transaction throughput than PostgreSQL.
-We attribute this performance gain to the proposed algorithm’s ability to schedule transactions more effectively than serializable snapshot-isolation Ports and Grittner (2012), a static concurrency control algorithm employed in PostgreSQL.
+We attribute this performance gain to the proposed algorithm’s ability to schedule transactions more effectively than serializable snapshot isolation Ports and Grittner (2012), a static concurrency control algorithm employed in PostgreSQL.
 We further evaluate the adaptability of our proposed algorithm against Polyjuice Wang et al.
-(2021), the state-of-the-art learned algorithm.
+(2021), state-of-the-art learned algorithm.
 Due to the limitation of constraining the transaction execution workflow, implementing Polyjuice on the codebase of NeurDB would be cumbersome.
 We therefore opt to implement our algorithm, named NeurDB(CC), into the Polyjuice codebase to facilitate a fair comparison.
-We set up a drift workload based on TPCC Council (2007) by varying the number of warehouses and threads.
-As shown in Figure 7(b), NeurDB(CC) adapts quickly to workload drifts and outperforms Polyjuice by up to 2.05$\times$.
+We set up a drift workload based on TPCC TPCC (2024) by varying the number of warehouses and threads.
+As shown in Figure 7(b), NeurDB(CC) adapts quickly to workload drift and outperforms Polyjuice by up to 2.05$\times$.
 The superior performance of NeurDB(CC) mainly stems from its design that encapsulates a fast yet accurate model to find the best concurrency control action, while facilitating the fine-tuning process with the two-phase adaptation algorithm.
@@
 We use stable models of Bao and Lero for the experiment, as they demonstrated good performance in their respective papers.
-As can be observed in Figure 8, NeurDB achieves up to 20.32% lower average latency of all evaluated queries, which demonstrates its effective adaptability to both data and workload drifts.
+As can be observed in Figure 8, NeurDB achieves up to 20.32% lower average latency of all evaluated queries, which demonstrates its effective adaptability to both data and workload drift.
 Due to the proposed dual-module model, NeurDB is able to effectively capture system conditions and use them to select an efficient query plan.
@@
 At that time, the fusion was far from reality due to limited development in both realms.
-With advancements in both fields over the years, numerous attempts by academia  Kraska et al.
+With advancements in AI and database fields over the years, numerous attempts by academia Kraska et al.
 (2019); Pavlo et al.
@@
 (2020); Li et al.
-(2021) have been made.
-However, many problems remain open, and it is important to consider database dynamicity in the fusion of AI with databases.
-Further, the system architecture may require a re-design to support AI-powered operations.
-NeurDB has been designed along this direction by seamlessly integrating AI into the database architecture, enabling continuous adaptation to data and workload drift.
+(2021) have been made to advance in-database AI analytics and autonomous DBMS optimization.
+However, many problems remain open, and it is important to consider data and workload drift in both directions.
+Here we shall briefly discuss recent advancements in AI and DBMS integration, with a particular focus on their approaches to handling data and workload drift.
+In-database AI Analytics
+Recent research increasingly focuses on optimizing the performance of in-database AI analytics.
+For example, some studies Salazar-Díaz et al.
+(2024); Park et al.
+(2022); Karanasos et al.
+(2020); Jasny et al.
+(2020) extend the existing query and transaction execution framework to improve the efficiency of analytics tasks.
+Further, in-database model management systems Su et al.
+(2024); Vartak (2017); Zhou et al.
+(2022) are proposed to minimize model storage costs while efficiently serving models for analytics tasks.
+These optimizations can be broadly adopted by full-fledged DBMSs supporting AI analytics, such as nsDB Yuan et al.
+(2024), GaussML Li et al.
+(2024), and NeurDB.
+However, apart from the mentioned approaches, NeurDB establishes an in-database AI ecosystem, specifically designed to enhance the efficiency and effectiveness of AI analytics under continuous data and workload drift.
+Autonomous DBMS Optimization
+Traditional AI-driven DBMS optimization in areas such as knob tuning Pavlo et al.
+(2017), resource management Li et al.
+(2023b, 2021), and learned system components Kraska et al.
+(2019); Marcus et al.
+(2021); Zhu et al.
+(2023); Wang et al.
+(2021); Kraska et al.
+(2018) typically relies on complete retraining to handle data and workload drift.
+Recently, the development of drift-handling mechanisms become popular, particularly for learned query optimizers and learned indexes.
+For example, some works introduce methods to automatically detect drift and trigger model updates Wu and Ives (2024); Li et al.
+(2022); Kurmanji and Triantafillou (2023), enabling learned query optimizers to withstand data and workload drift.
+Further, updatable learned indexes Zhang et al.
+(2024); Ding et al.
+(2020); Wu et al.
+(2021) are enhanced with the adaptability to continuous transactional updates.
+Unlike existing works focusing on certain components, NeurDB introduces a unified architecture based on the proposed in-database AI ecosystem to facilitate fast adaptation across all its key components.
+We are extending NeurDB to be an intelligent cloud-native HTAP database with built-in support for advanced AI analytics.
+In particular, we disaggregate the system architecture into separate compute and storage layers Pang and Wang (2024); Cao et al.
+(2021), enabling dedicated compute nodes to handle OLTP, OLAP, and AI analytics independently.
+Based on this foundation, we are devising AI-powered resource scheduling to enable on-demand resource allocation and component assembling, thus enhancing system scalability and adaptability.
 7. Conclusions
 This paper presents NeurDB, a novel AI-powered autonomous DBMS that is adaptable to data and workload drift.
-NeurDB, with its fast-adaptive learned system components and in-database AI ecosystem, facilitates efficient and effective in-database AI analytics.
+NeurDB, with its fast-adaptive learned database components and in-database AI ecosystem, facilitates efficient and effective in-database AI analytics and autonomous system optimization.
 Empirical evaluations demonstrate the superiority of NeurDB, highlighting its potential to realize a seamless fusion of AI and databases.
+{acks}
+We would like to thank Jiaqi Zhu for her help and comments.
+The work of NUS researchers is partially supported by the Lee Foundation in terms of Beng Chin Ooi’s Lee Kong Chian Centennial Professorship fund and NUS Faculty Development Fund.
+Gang Chen’s work is supported by National Key Research and Development Program of China (2022YFB2703100).
+Yanyan Shen’s work is supported by National Key Research and Development Program of China (2022YFE0200500).
+Yuncheng Wu’s work is supported by National Key Research and Development Program of China (2023YFB4503600).
+Meihui Zhang’s work is supported by National Natural Science Foundation of China (62072033).